OpenStudy (anonymous):

What is the solution to r over three is equal to one over three? one over twelve twelve over one three over four four over three

OpenStudy (anonymous):

Where's the option for none of the above?

OpenStudy (anonymous):

r/3 = 1/3 by observation r must be = 1